# Asymptotic Analysis of High Order IMEX-RK Methods for ES-BGK Model at Navier-Stokes level

### v2

@@ 3 Asymptotic Properties of the IMEX RK Schemes
and $V^{n}$, $A(V^{n})$, $B(V^{n})$ are given by (3.57), asymptotically becomes consistent time discretization of the CNS equations (1.21).
Proof.
The IMEX RK scheme of type II applied to (1.7) can be written in compact form as
@@
[…]
into (3.75a), multiplying by $\phi(v)$ function and integrating on $v$, we get
$\begin{split}U^{(1)}&=U^{n},\\[5.69054pt] \hat{\bf U}&=U^{n}\hat{{\bf e}}-\Delta t(\tilde{a}\nabla_{x}F(U^{n})+\hat{\tilde{A}}\nabla_{x}F(\hat{\bf U}]))-\varepsilon\Delta t\left(\tilde{a}\langle\phi L(f_{1}^{n})\rangle+\hat{\tilde{A}}\langle\phi L(\hat{\bf f}_{1})\rangle\right),\\[5.69054pt] U^{n+1}&=U^{n}-\Delta t(\tilde{b}_{1}\nabla_{x}F(U^{n})+\hat{\tilde{b}}^{\top}\nabla_{x}F(\hat{\bf U}))-\varepsilon\Delta t\left(\tilde{b}_{1}\langle\phi L(f_{1}^{n})\rangle+\hat{\tilde{b}}^{\top}\langle\phi L(\hat{\bf f}_{1})\rangle\right),\end{split}$ (3.77)
@@
Now we evaluate $f_{1}^{n}$ ad $\hat{\bf f}_{1}$ in (3.77). From (3.75a) and substituting (3.76) we get
$\begin{split}\mathcal{M}[\hat{\bf F}]=\displaystyle\mathcal{M}[f^{n}]\hat{{\bf e}}&-\Delta t\left(\tilde{a}L(\mathcal{M}[f^{n}])+\hat{\tilde{A}}\,L(\mathcal{M}[\hat{\bf F}])\right)\\ &-\varepsilon\left({\hat{\bf f}}_{1}-f_{1}^{n}{\bf e}+\Delta t\left(\tilde{a}L(f_{1}^{n})+\hat{\tilde{A}}\,L(\hat{\bf f}_{1})\right)\right)+\\ &\Delta t\left(a\tau^{n}(g^{n}-f_{1}^{n})+\hat{A}\hat{\bar{\tau}}(\hat{\bf g}-\hat{{\bf f}}_{1})\right).\end{split}$ (3.79)
Following the same argument for type I in (3.56)-(3.59), we obtain
@@
where ${\bf u}^{\top}=(u^{(1)},\hat{\bf u}^{\top})$, ${\bf T}^{\top}=(T^{(1)},...,\hat{\bf T}^{\top})$ with $\hat{\bf u}^{\top}\in\mathbb{R}^{d_{v}\times(s-1)}$, $\hat{\bf T}^{\top}\in\mathbb{R}^{s-1}$ and
@@
$\tau^{n}(f_{1}^{n}-g^{n})=-\mathcal{M}[f^{n}]\left(A(V^{n}):\frac{\sigma(u^{n})}{2}+2B(V^{n})\cdot\nabla_{x}\sqrt{T^{n}}\right)+\mathcal{O}(\varepsilon).$
and substituting it into (3.80), we get
Inserting (3.74) and (3.81) into (3.77), we get
$\begin{split}U^{1}&=U^{n},\\ \hat{\bf U}&=U^{n}\hat{\bf e}-\Delta t\left(\tilde{a}\nabla_{x}F(U^{n})+\hat{\tilde{A}}\nabla_{x}F(\hat{\bf U})\right)-\varepsilon\Delta t(\tau^{n})^{-1}\hat{\tilde{A}}\hat{A}^{-1}\left(a-\tilde{a}\right)\nabla_{x}\mathcal{F}(U^{n})\\ &-\varepsilon\Delta t\tilde{a}\nabla_{x}\left(G(U^{n})-(\tau^{n})^{-1}\mathcal{F}(U^{n})\right)-\varepsilon\Delta t\hat{\tilde{A}}\nabla_{x}\left(\mathcal{G}(\hat{\bf U})-(\hat{\bar{\tau}})^{-1}\hat{A}^{-1}\hat{\tilde{A}}\mathcal{F}(\hat{\bf U})\right)\\ &+\mathcal{O}(\varepsilon^{2}),\end{split}$ (3.82)
with the numerical solution
$\begin{split}U^{n+1}&=U^{n}-\Delta t\left(\tilde{b}_{1}\nabla_{x}F(U^{n})+\hat{\tilde{b}}^{\top}\nabla_{x}F({\bf U})\right)-\varepsilon\Delta t(\tau^{n})^{-1}\hat{\tilde{b}}^{\top}\hat{A}^{-1}(\tilde{a}-a)\nabla_{x}\mathcal{F}(U^{n})\\ &-\varepsilon\Delta t\tilde{b}_{1}\nabla_{x}\left(G(U^{n})-(\tau^{n})^{-1}\mathcal{F}(U^{n})\right)-\varepsilon\Delta t\hat{\tilde{b}}^{\top}\nabla_{x}\left(\mathcal{G}(\hat{\bf U})-\hat{\tau}^{-1}\hat{A}^{-1}\hat{\tilde{A}}\mathcal{F}({\bf U})\right)\\ &+\mathcal{O}(\varepsilon^{2}).\end{split}$ (3.83)
Note that (3.82) and (3.83) can be simplified by the condition (3.73). Now comparing (3.82) and (3.83) with (3.78) we get
$\begin{split}H(U^{n})&=G(U^{n})-(\tau_{n})^{-1}\mathcal{F}(U^{n}){+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right)},\\ H(\hat{\bf U})&=G(\hat{\bf U})-\hat{\tau}^{-1}\hat{A}^{-1}\hat{\tilde{A}}\mathcal{F}(\hat{\bf U}){+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right)}.\end{split}$ (3.84)
Then, from (3.84) we see that the viscosity and thermal conductivity is given by
$\mu^{(1)}=\frac{1}{(1-\nu)\tau^{n}}\,p^{n}\textrm{b},\quad\kappa^{(1)}=\frac{d_{v}+2}{2\tau^{n}}{p}^{n}\textrm{b},$ (3.85)
with $\textrm{b}=\hat{A}^{-1}\tilde{a}$, and $s-1\times s-1$ matrices ${\hat{\bar{\mu}}}=(\mu_{ij})$ and ${\hat{\bar{\kappa}}}=(\kappa_{ij})$
$\hat{{\bar{\mu}}}=\frac{1}{(1-\nu)}\hat{\tau}^{-1}(\hat{A}^{-1}\hat{\tilde{A}})\,\hat{\bar{p}}e^{\top},\quad\hat{{\bar{\kappa}}}=\frac{d_{v}+2}{2}\,\hat{\tau}^{-1}(\hat{A}^{-1}\hat{\tilde{A}})\hat{\bar{p}}e^{\top},$ (3.86)
with $\hat{\bar{p}}=(p^{(2)},...,p^{(s)})^{\top}$. Therefore from (3.78) we obtain
$\begin{split}U^{(1)}&=U^{n},\\ \hat{\bf U}&=U^{n}{\bf e}-\Delta t\tilde{a}\nabla_{x}F(U^{n})-\Delta t\tilde{A}\nabla_{x}F(\hat{\bf U})+\varepsilon\Delta t\tilde{a}S(U^{n})+\varepsilon\Delta t\hat{\tilde{A}}S(\hat{\bf U})+\mathcal{O}(\varepsilon^{2})\\[5.69054pt] U^{n+1}&=U^{n}-\Delta t\tilde{b}_{1}\nabla_{x}F(U^{n})-\Delta t\hat{\tilde{b}}^{\top}\nabla_{x}F(\hat{\bf U})+\varepsilon\Delta t\tilde{b}_{1}S(U^{n})+\varepsilon\Delta t\hat{\tilde{b}}^{\top}S(\hat{\bf U})+\mathcal{O}(\varepsilon^{2})\end{split}$ (3.87)
@@
where the viscosity and thermal conductivity are given by (3.85) and (3.86). The equations (3.87) can be written as
${\bf U}=U^{n}{\bf e}-\Delta t\tilde{A}\nabla_{x}F({\bf U})+\varepsilon\Delta t{\tilde{A}}S({\bf U})+\mathcal{O}(\varepsilon^{2})$ (3.88)
$U^{n+1}=U^{n}-\Delta t\tilde{b}^{\top}\nabla_{x}F({\bf U})+\varepsilon\Delta t\tilde{b}^{\top}S({\bf U})+\mathcal{O}(\varepsilon^{2})$ (3.89)

### v3

@@ 3 Asymptotic Properties of the IMEX RK Schemes
As in the case of Theorem 3.6, we remark that the order $p$ could be ensured by enforcing additional order conditions derived up to order $3$ for type II from the associated two pair of Butcher tableaux in [8].
Proof.
The IMEX RK scheme of type II applied to (1.7) can be written in compact form (2.30) as
@@
Now inserting expansions
into (3.71), multiplying by $\phi(v)$ function and integrating on $v$, we get
$\begin{split}U^{(1)}&=U^{n},\\[5.69054pt] \hat{\bf U}&=\hat{{\bf e}}\otimes_{K}U^{n}-\Delta t\left(\tilde{a}\otimes_{K}\nabla_{x}\cdot F(U^{n})+{\bf\hat{\tilde{A}}}\nabla_{x}\cdot F(\hat{\bf U})\right)-\varepsilon\Delta t\left(\tilde{a}\otimes_{K}\langle\phi L(f_{1}^{n})\rangle+{\bf\hat{\tilde{A}}}\langle\phi L(\hat{\bf f}_{1})\rangle\right),\\[5.69054pt] U^{n+1}&=U^{n}-\Delta t\left(\tilde{b}_{1}\nabla_{x}\cdot F(U^{n})+{\bf\hat{\tilde{b}}}\nabla_{x}\cdot F(\hat{\bf U})\right)-\varepsilon\Delta t\left(\tilde{b}_{1}\langle\phi L(f_{1}^{n})\rangle+{\bf\hat{\tilde{b}}}\langle\phi L(\hat{\bf f}_{1})\rangle\right).\end{split}$ (3.73)
[…]
$\begin{split}\hat{\bf U}&=\begin{pmatrix}U^{(2)},U^{(3)},\cdots,U^{(s)}\end{pmatrix}^{\top}\in\mathbb{R}^{(2+d_{v})(s-1)},\\ U^{(i)}&=(\rho^{(i)},\rho^{(i)}u^{(i)},E^{(i)})^{\top}\in\mathbb{R}^{2+d_{v}}\\ \nabla_{x}\cdot F(\hat{\bf U})&=\left(\nabla_{x}\cdot F(U^{(2)}),\nabla_{x}\cdot F(U^{(3)}),...,\nabla_{x}\cdot F(U^{(s)})\right)^{\top}\in\mathbb{R}^{(2+d_{v})(s-1)},\\ \nabla_{x}\cdot F(U^{(i)})&=\left(\begin{array}[]{c}\nabla_{x}\cdot(\rho^{(i)}u^{(i)})\\ \nabla_{x}\cdot(\rho^{(i)}u^{(i)}\otimes u^{(i)}+\rho^{(i)}T^{(i)}Id)\\ \nabla_{x}\cdot((E^{(i)}+\rho^{(i)}T^{(i)})u^{(i)})\end{array}\right)\in\mathbb{R}^{2+d_{v}}\,(i=1,...,s).\end{split}$ (3.74)
Then, by Lemma A.3 we obtain
[…]
where
[…]
$\nabla_{x}\cdot H(U^{(i)})$ $=\left(\begin{array}[]{c}0\\ \nabla_{x}\cdot(\rho^{(i)}\Theta_{1}^{(i)})\\ \nabla_{x}\cdot({\mathbb{Q}_{1}^{(i)}}+\rho^{(i)}{\Theta_{1}^{(i)}}{u^{(i)}})\end{array}\right)\in\mathbb{R}^{2+d_{v}},\quad i=1,...,s.$
Now we evaluate $\hat{\bf f}_{1}$ in (3.73). Substituting (3.72) into (3.71), we obtain
$\begin{split}\mathcal{M}[\hat{\bf F}]=\displaystyle\mathcal{M}[f^{n}]\hat{{\bf e}}&-\Delta t\left(\tilde{a}L(\mathcal{M}[f^{n}])+\hat{\tilde{A}}\,L(\mathcal{M}[\hat{\bf F}])\right)\\ &-\varepsilon\left({\hat{\bf f}}_{1}-f_{1}^{n}{\bf e}+\Delta t\left(\tilde{a}L(f_{1}^{n})+\hat{\tilde{A}}\,L(\hat{\bf f}_{1})\right)\right)\\ &+\Delta t\left(a\tau^{n}(g^{n}-f_{1}^{n})+\hat{A}\hat{\bar{\tau}}(\hat{\bf g}-\hat{{\bf f}}_{1})\right).\end{split}$ (3.76)
Following the same argument for type I in (3.57)-(3.59), we have
@@
where ${\bf u}^{\top}=(u^{(1)},\hat{\bf u}^{\top})$, ${\bf T}^{\top}=(T^{(1)},\hat{\bf T}^{\top})$ with $\hat{\bf u}^{\top}\in\mathbb{R}^{d_{v}(s-1)}$, $\hat{\bf T}^{\top}\in\mathbb{R}^{s-1}$ and
$\displaystyle{\textbf{V}}^{\top}=(V^{(1)},\hat{V}^{\top}),\quad{\hat{\bf V}}=(V^{(2)},V^{(2)},...,V^{(s)})^{\top}\in\mathbb{R}^{d_{v}(s-1)}\quad V^{(i)}=\frac{v-u^{(i)}}{\sqrt{T^{(i)}}}\in\mathbb{R}^{d_{v}},$
$\tau^{n}(f_{1}^{n}-g^{n})=-\mathcal{M}[f^{n}]\left(A(V^{n}):\frac{\sigma(u^{n})}{2}+2B(V^{n})\cdot\nabla_{x}\sqrt{T^{n}}\right)+\mathcal{O}(\varepsilon),$
and substituting this into (3.77), we get
Recall that $F^{(1)}=f^{n}$, $U^{(1)}=U^{n}$, we set $g^{(1)}=g^{n}$, $f_{1}^{(1)}=f_{1}^{n}$. Now, we multiplying by $v\otimes v$ or $v|v|^{2}/2$ both side in (3.78), and take the integration over $v$, which together with (A.113), (A.114), (A.115) gives
$\begin{split}{\rho^{(1)}\Theta_{1}^{(1)}}&=\nu{\rho^{(1)}\Theta_{1}^{(1)}}-\frac{1}{\tau^{n}}\rho^{n}T^{n}\sigma(u^{n})+\mathcal{O}\left({\varepsilon}\right),\\ \mathbb{Q}_{1}^{(1)}+\rho^{(1)}\Theta_{1}^{(1)}u^{(1)}&=\nu{\rho^{(1)}\Theta_{1}^{(1)}}u^{(1)}-{\frac{1}{\tau^{n}}}\left(\rho^{n}T^{n}\sigma(u^{n})u^{n}+\frac{d_{v}+2}{2}\rho^{n}T^{n}\nabla_{x}T^{n}\right)+\mathcal{O}\left({\varepsilon}\right),\end{split}$ (3.79)
and
$\begin{split}{\rho^{(i)}\Theta_{1}^{(i)}}&=\nu{\rho^{(i)}\Theta_{1}^{(i)}}-\frac{1}{\tau^{(i)}}(\hat{A}^{-1}(\tilde{a}-a))_{i-1}\rho^{n}T^{n}\sigma(u^{n})\\ &-{{\frac{1}{\tau^{(i)}}\sum_{j=1}^{i-2}(\hat{A}^{-1}\hat{\tilde{A}})_{ij}}}\rho^{(j+1)}T^{(j+1)}\sigma(u^{(j+1)})+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right),\\ \mathbb{Q}_{1}^{(i)}+\rho^{(i)}\Theta_{1}^{(i)}u^{(i)}&=\nu{\rho^{(i)}\Theta_{1}^{(i)}}u^{(i)}-{\frac{1}{\tau^{(i)}}(\hat{A}^{-1}(\tilde{a}-a))_{i-1}}\left(\rho^{n}T^{n}\sigma(u^{n})u^{n}+\frac{d_{v}+2}{2}\rho^{n}T^{n}\nabla_{x}T^{n}\right)\\ &-{\frac{1}{\tau^{(i)}}\sum_{j=1}^{i-2}(\hat{A}^{-1}\hat{\tilde{A}})_{ij}}\left(\rho^{(j+1)}T^{(j+1)}\sigma(u^{(j+1)})u^{(j+1)}+\frac{d_{v}+2}{2}\rho^{(j+1)}T^{(j+1)}\nabla_{x}T^{(j+1)}\right)+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right),\quad i=2,...,s\end{split}$
with $\hat{A}^{-1}(\tilde{a}-a)\in\mathbb{R}^{s-1}$. Thus, the approximations of stress tensor and heat flux are given by
$\rho^{(1)}{\Theta_{1}^{(1)}}$ $=\frac{1}{(1-\nu)\tau^{n}}\rho^{n}T^{n}\sigma(u^{n})+\mathcal{O}\left({\varepsilon}\right),\quad\mathbb{Q}_{1}^{(1)}=\frac{d_{v}+2}{2\tau^{n}}\rho^{n}T^{n}\nabla_{x}T^{n}+\mathcal{O}\left({\varepsilon}\right),$
$\rho{\bf\hat{\Theta}_{1}}$ $=\begin{pmatrix}\rho^{(2)}{\Theta_{1}^{(2)}},\rho^{(3)}{\Theta_{1}^{(3)}},\cdots,\rho^{(s)}{\Theta_{1}^{(s)}}\end{pmatrix}^{\top}$,
$=-\frac{1}{1-\nu}\left(\hat{\bar{\tau}}^{-1}\hat{A}^{-1}(\tilde{a}-a)p^{n}\right)\otimes_{K}\sigma(U^{n})-\frac{1}{1-\nu}\left(\hat{\bar{\tau}}^{-1}\hat{A}^{-1}\hat{\tilde{A}}\,diag(\hat{\bar{p}})\otimes_{K}Id\right)\times\sigma(\hat{\bf u})+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right)$
$=-\bar{\mu}^{(1)}\otimes_{K}\sigma(U^{n})-\left(\hat{\bar{\mu}}\otimes_{K}Id\right)\times\sigma({\hat{\bf u}})+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right)$
${\bf q}=\mathbb{Q}_{1}$ $=\begin{pmatrix}\mathbb{Q}_{1}^{(2)},\mathbb{Q}_{1}^{(3)},\cdots,\mathbb{Q}_{1}^{(s)}\end{pmatrix}^{\top}$,
$=-\frac{d_{v}+2}{2}\left({\hat{\bar{\tau}}}^{-1}{\hat{A}}^{-1}(\tilde{a}-a)p^{n}\right)\otimes_{K}\nabla_{x}{T^{n}}-\frac{d_{v}+2}{2}\left({\hat{\bar{\tau}}}^{-1}{\hat{A}}^{-1}\hat{\tilde{A}}\,diag(\hat{\bar{p}})\otimes_{K}Id\right)\times\nabla_{x}\hat{{\bf T}}+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right)$
$=-\bar{\kappa}^{(1)}\otimes_{K}\nabla_{x}{T^{n}}-(\hat{\bar{\kappa}}\otimes_{K}Id)\times\nabla_{x}\hat{{\bf T}}+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right)$
where $\sigma(\hat{{\bf u}})=(\sigma(u^{(2)}),\,...,\,\sigma(u^{(s)}))^{\top}$, $\nabla_{x}\hat{{\bf T}}=(\nabla_{x}T^{(2)},\,...,\,\nabla_{x}T^{(s)})^{\top}$, $\sigma(u^{(i)})=\nabla_{x}u^{(i)}+(\nabla_{x}u^{(i)})^{\top}-\frac{2}{d_{v}}\nabla_{x}\cdot u^{(i)}Id$ (i=2,,si=2,...,s) and $\hat{\bar{p}}=(p^{(2)},p^{(3)},...,p^{(s)})^{\top}$ with $p^{(i)}=\rho^{(i)}T^{(i)}$ (i=2,,si=2,...,s). Now we introduce the viscosity and thermal conductivity matrices given by
@@
where
$\bar{\mu}^{(1)}=\frac{1}{1-\nu}\left(\hat{\bar{\tau}}^{-1}\hat{A}^{-1}(\tilde{a}-a)p^{n}\right),\quad\bar{\kappa}^{(1)}=\frac{d_{v}+2}{2}\left({\hat{\bar{\tau}}}^{-1}{\hat{A}}^{-1}(\tilde{a}-a)p^{n}\right),$ (3.81)
and ${\hat{\bar{\mu}}}=(\mu_{ij})$ and ${\hat{\bar{\kappa}}}=(\kappa_{ij})$ are $s-1\times s-1$ matrices such that
[…]
Therefore, from (3.75) we obtain
$\begin{split}U^{(1)}&=U^{n},\\ \hat{\bf U}&=\hat{\bf e}\otimes_{K}U^{n}-\Delta t\tilde{a}\otimes_{K}\nabla_{x}\cdot F(U^{n})-\Delta t{\bf\hat{\tilde{A}}}\nabla_{x}\cdot F(\hat{\bf U})+\varepsilon\Delta t\tilde{a}\otimes_{K}\nabla\cdot S(U^{n})+\varepsilon\Delta t{\bf\hat{\tilde{A}}}\nabla_{x}\cdot\overline{S}(\hat{\bf U})+\mathcal{O}(\varepsilon^{2})\\[5.69054pt] U^{n+1}&=U^{n}-\Delta t\tilde{b}_{1}\nabla_{x}\cdot F(U^{n})-\Delta t{\bf\hat{\tilde{b}}}\nabla_{x}\cdot F(\hat{\bf U})+\varepsilon\Delta t\tilde{b}_{1}\nabla_{x}\cdot S(U^{n})+\varepsilon\Delta t{\bf\hat{\tilde{b}}}\nabla_{x}\cdot\overline{S}(\hat{\bf U})+\mathcal{O}(\varepsilon^{2})\end{split}$ (3.83)
@@
$\begin{split}\nabla_{x}\cdot S(U^{n})&=\left(\begin{array}[]{c}\nabla_{x}\cdot{\bf 0},\\ \nabla_{x}\cdot({\mu^{n}}\sigma(u^{n})),\\ \nabla_{x}\cdot({\mu^{n}}\sigma(u^{n})u^{n}+{\kappa^{n}}\nabla_{x}T^{n})\end{array}\right),\quad\mu^{n}=\frac{1}{(1-\nu)\tau^{n}}p^{n},\quad\kappa^{n}=\frac{d_{v}+2}{2\tau^{n}}p^{n},\cr\nabla_{x}\cdot\overline{S}(\hat{{\bf U}})&=\left(\nabla_{x}\cdot\overline{S}(U^{(2)}),\nabla_{x}\cdot\overline{S}(U^{(3)}),...,\nabla_{x}\cdot\overline{S}(U^{(s)})\right)^{\top}\in\mathbb{R}^{(d_{v}+2)(s-1)},\\ \nabla_{x}\cdot\overline{S}(U^{(i)})&=\left(\begin{array}[]{c}\nabla_{x}\cdot{\bf 0},\\ \nabla_{x}\cdot\left({\bar{\mu}^{(1)}_{i-1}}\sigma(u^{n})\right)+\nabla_{x}\cdot\left(\sum_{j=1}^{i-2}\hat{\bar{\mu}}_{ij}\sigma(u^{(j+1)})\right)\\ \nabla_{x}\cdot\left(\bar{\mu}^{(1)}_{i-1}\sigma(u^{n})u^{n}+{\bar{\kappa}^{(1)}_{i-1}\nabla_{x}T^{n}}\right)+\nabla_{x}\cdot\left(\sum_{j=1}^{i-2}\hat{\bar{\mu}}_{ij}\sigma(u^{(j+1)})u^{(j)}+{\hat{\bar{\kappa}}_{ij}\nabla_{x}T^{(j+1)}}\right)\end{array}\right)\in\mathbb{R}^{d_{v}+2}.\end{split}$
From $\tau^{(i)}=\tau^{(j)}+\mathcal{O}(\Delta t)+\mathcal{O}(\varepsilon^{2})$ for i,j=1,,si,j=1,...,s, we further get
$\begin{split}U^{(1)}&=U^{n},\\ \hat{\bf U}&=\hat{\bf e}\otimes_{K}U^{n}-\Delta t\tilde{a}\otimes_{K}\nabla_{x}\cdot F(U^{n})-\Delta t{\bf\hat{\tilde{A}}}\nabla_{x}\cdot F(\hat{\bf U})\\ &+\varepsilon\Delta t\tilde{a}\otimes_{K}\nabla\cdot S(U^{n})+\varepsilon\Delta t{\bf\hat{\tilde{A}}}\nabla_{x}\cdot\overline{\overline{S}}(\hat{\bf U})+\mathcal{O}(\varepsilon^{2})+\mathcal{O}(\varepsilon\Delta t^{2})\\[5.69054pt] U^{n+1}&=U^{n}-\Delta t\tilde{b}_{1}\nabla_{x}\cdot F(U^{n})-\Delta t{\bf\hat{\tilde{b}}}\nabla_{x}\cdot F(\hat{\bf U})\\ &+\varepsilon\Delta t\tilde{b}_{1}\nabla_{x}\cdot S(U^{n})+\varepsilon\Delta t{\bf\hat{\tilde{b}}}\nabla_{x}\cdot\overline{\overline{S}}(\hat{\bf U})+\mathcal{O}(\varepsilon^{2})+\mathcal{O}(\varepsilon\Delta t^{2})\end{split}$
with
$\begin{split}\nabla_{x}\cdot\overline{\overline{S}}({\bf U})&=\left(\nabla_{x}\cdot\overline{\overline{S}}(U^{(2)}),\nabla_{x}\cdot\overline{\overline{S}}(U^{(3)}),...,\nabla_{x}\cdot\overline{\overline{S}}(U^{(s)})\right)^{\top}\in\mathbb{R}^{(d_{v}+2)(s-1)},\\ \nabla_{x}\cdot\overline{\overline{S}}(U^{(i)})&=\left(\begin{array}[]{c}\nabla_{x}\cdot{\bf 0}\\ \left(\hat{A}^{-1}(\tilde{a}-a)\right)_{i-1}\nabla_{x}\cdot\left(\mu^{n}\sigma(u^{n})\right)\\ \left(\hat{A}^{-1}(\tilde{a}-a)\right)_{i-1}\nabla_{x}\cdot\left(\left(\mu^{n}\sigma(u^{n})u^{n}\right)+\kappa^{n}\nabla_{x}T^{n}\right))\end{array}\right)\\ &+\left(\begin{array}[]{c}\nabla_{x}\cdot{\bf 0}\\ \sum_{j=1}^{i-2}(\hat{A}^{-1}\hat{\tilde{A}})_{ij}\nabla_{x}\cdot\left(\mu^{(j+1)}\sigma(u^{(j+1)})\right)\\ \sum_{j=1}^{i-2}(\hat{A}^{-1}\hat{\tilde{A}})_{ij}\nabla_{x}\cdot\left(\mu^{(j+1)}\sigma(u^{(j+1)})u^{(j+1)}+\kappa^{(j+1)}\nabla_{x}T^{(j+1)}\right)\end{array}\right)\end{split}$
where $\mu^{(j)}=\frac{1}{(1-\nu)\tau^{(j)}}p^{(j)}$, $\kappa^{(j)}=\frac{d_{v}+2}{2\tau^{(j)}}p^{(j)}$. Defining
$\begin{split}\nabla_{x}\cdot S(\hat{{\bf U}})&=\left(\nabla_{x}\cdot S(U^{(2)}),\nabla_{x}\cdot S(U^{(3)}),...,\nabla_{x}\cdot S(U^{(s)})\right)^{\top}\in\mathbb{R}^{(d_{v}+2)(s-1)},\\ \nabla_{x}\cdot S(U^{(i)})&=\left(\begin{array}[]{c}\nabla_{x}\cdot{\bf 0},\\ \nabla_{x}\cdot\left(\mu^{(i)}\sigma(u^{(i)})\right),\\ \nabla_{x}\cdot\left(\mu^{(i)}\sigma(u^{(i)})u^{(i)}+\kappa^{(i)}\nabla_{x}T^{(i)}\right)\end{array}\right)\in\mathbb{R}^{d_{v}+2},\quad i=2,...,s,\end{split}$
we finally obtain
$\begin{split}U^{(1)}&=U^{n},\\ \hat{\bf U}&=\hat{\bf e}\otimes_{K}U^{n}-\Delta t\tilde{a}\otimes_{K}\nabla_{x}\cdot F(U^{n})-\Delta t{\bf\hat{\tilde{A}}}\nabla_{x}\cdot F(\hat{\bf U})\\ &+\varepsilon\Delta t\tilde{a}\otimes_{K}\nabla_{x}\cdot S(U^{n})+\varepsilon\Delta t{\hat{\tilde{A}}\hat{A}^{-1}(\tilde{a}-a)}\otimes_{K}\nabla_{x}\cdot S(U^{n})+\varepsilon\Delta t{\bf\hat{B}}\nabla_{x}\cdot S(\hat{\bf U})+\mathcal{O}(\varepsilon^{2})+\mathcal{O}(\varepsilon\Delta t^{2})\\ U^{n+1}&=U^{n}-\Delta t\tilde{b}_{1}\nabla_{x}\cdot F(U^{n})-\Delta t{\bf\hat{\tilde{b}}}\nabla_{x}\cdot F(\hat{\bf U})\cr&+\varepsilon\Delta t\tilde{b}_{1}\nabla_{x}\cdot S(U^{n})+\varepsilon\Delta t{\hat{\tilde{b}}^{\top}\hat{A}^{-1}(\tilde{a}-a)}\nabla_{x}\cdot S(U^{n})+\varepsilon\Delta t\boldsymbol{\omega}\nabla_{x}\cdot S(\hat{\bf U})+\mathcal{O}(\varepsilon^{2})+\mathcal{O}(\varepsilon\Delta t^{2})\end{split}$ (3.84)
where ${\bf\hat{B}}=\hat{B}\otimes_{K}I_{d_{v}+2}$ with $\hat{B}=\hat{\tilde{A}}\hat{A}^{-1}\hat{\tilde{A}}$ and $\hat{\boldsymbol{\omega}}=\hat{\omega}\otimes_{K}I_{d_{v}+2}$ with $\hat{\omega}=\hat{\tilde{b}}^{\top}\hat{A}^{-1}\hat{\tilde{A}}$. By the extra condition (3.68), we get
$\begin{split}U^{(1)}&=U^{n},\\ \hat{\bf U}&=\hat{\bf e}\otimes_{K}U^{n}-\Delta t\tilde{a}\otimes_{K}\nabla_{x}\cdot F(U^{n})-\Delta t{\bf\hat{\tilde{A}}}\nabla_{x}\cdot F(\hat{\bf U})\\ &+\varepsilon\Delta t{\mathbb{\bf b}_{1}}\otimes_{K}\nabla_{x}\cdot S(U^{n})+\varepsilon\Delta t{\bf\hat{B}}\nabla_{x}\cdot S(\hat{\bf U})+\mathcal{O}(\varepsilon^{2})+\mathcal{O}(\varepsilon\Delta t^{2})\\ U^{n+1}&=U^{n}-\Delta t\tilde{b}_{1}\nabla_{x}\cdot F(U^{n})-\Delta t{\bf\hat{\tilde{b}}}\nabla_{x}\cdot F(\hat{\bf U})\cr&+\varepsilon\Delta t\omega_{1}\nabla_{x}\cdot S(U^{n})+\varepsilon\Delta t\hat{\boldsymbol{\omega}}\nabla_{x}\cdot S(\hat{\bf U})+\mathcal{O}(\varepsilon^{2})+\mathcal{O}(\varepsilon\Delta t^{2})\end{split}$
where and $\mathbb{\bf b}_{1}={\hat{\tilde{A}}\hat{A}^{-1}\tilde{a}}$ and $\omega_{1}={\hat{\tilde{b}}^{\top}\hat{A}^{-1}\tilde{a}}$. Next, we consider the following explicit-type RK method of order $p$ applied to the CNS equations (1.21):